OpenStudy (anonymous):

Find the general solution for the following differential equation: y' + 8y = -6y^-8 What substitution will you make?

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

@wio

OpenStudy (loser66):

I don't see any substitution needed there, just y' = \(\dfrac {-6}{y^8}-8y= -\dfrac{6+8y^9}{y^8}\) then \[\frac{dy}{dx}=-\frac{6+y^9}{y^8}\rightarrow \frac{y^8}{6+y^9}dy=-dx\] integral both sides, that's it

OpenStudy (anonymous):

thats what i got but its online and its asking for a sub

OpenStudy (anonymous):

the sub is v(t) =y^9 only thing i cant get is the general solution

OpenStudy (anonymous):

This looks like a Bernoulli equation, which would explain the choice of substitution.